Tohban Report 2012-11-07

From RHESSI Wiki

Jump to: navigation, search

Tohban Reports
Start Date: 31 Oct 2012
End Date: 7 Nov 2012
Tohban: Hazel Bain
Tohban email:
Next Tohban: Claire Raftery
List all reports


Solar Activity

Solar activity this week was low with only 2 C class flares. There are currently 8 active regions on disk, a couple of which are due to rotate off the disk within the next 24 hours.

How many GOES flares occurred?

 Flares above B, C, M, X class were     14     2     0     0

And how many of these are listed in the RHESSI flare list?

 Flares above B, C, M, X class were     11     1     0     0

And how many had EXCELLENT coverage?

 Flares above B, C, M, X class were     10     0     0     0

There were RHESSI flares/GOES flares 65 / 16 over the time range 31-Oct-12 07-Nov-12

Memory Management

The SSR reached a maximum of 65% during the week due to missed Wallops passes last as a result of Hurricane Sandy. Extra Weilheim passes were added to download the data. The SSR is now reaching ~5% at the end of the daily pass set.

Spacecraft Status

Cold plate 1 and 2 temperatures are 108.6 K and 106.3 K respectively.

Data Gaps

There were a number of small 5 minute data gaps between the 2nd and 4th of November.

Detector issues

Detector 8 continued to show an increases in noise coming out of SAA. The HV was lowered by 5 steps from 3456 V to 3358 V at 2012-310-19:31:18 and another 5 steps to 3260 V at 2012-310-21:10:52. However further action may still be required.

The front fast and slow thresholds on detector 4 were optimized during 2 passes on the 6th of November. We were able to reduce the front slow threshold significantly, now down to the nominal value used by segmented detectors, without seeing a large increase in noise. We were able to lower the front fast threshold a little as well, but not by nearly as much.

G4 Front Slow 0x70 to 0x0C.

G4 Front Fast 0x60 to 0x50.

We also noted that there was quite a bit of noise on detector 2, so we raised the front slow threshold on detector 2. This may help with the SSR fill level issues we have been seeing recently.

G2 Front Slow 0x60 to 0x78.

Other notes

The new tohban should keep an eye on the monitor rates for detectors 2, 4 and 8 for issues/noise that might arise as a result of this weeks operations. If raising the detector 2 slow threshold hels lower the SSR fill level then we should consider stepping the decimation level down from Active/vigorous to normal/vigorous.

Spacecraft Management

Decimation Active/vigorous
HLAT Decimation Rear decimation weight 6, no front decimation
Night time data (fronts) +/- 4 minutes
Night time data (rears) +/- 4 minutes
Require extra passes? Yes
Requirement for moving pointer? No
Attenuator operation Normal
Detector problems? Operations to det 2, 4 and 8, see above for details.
Personal tools